What is the probability for each sample that a rope is outside 2 standard deviations (above or below) the sample mean?

Mean Median Mode Standard Dev Coefficient of Variation/Covariance
Sample 1 83.77 84.05 83.88 83.94 84.01 84.03 84.06 84.07 84.09 84.11 84.001 84.04 #N/A 0.1070 0.12740972
Sample 2 84.01 84.09 83.98 83.95 83.92 83.85 83.82 83.83 83.87 83.88 83.92 83.9 #N/A 0.0873 0.10403401
Answer #1

for sample 1 ; 2 std deviation away values from mean =84.001-/+2*0.1070=83.7870 to 84.2150

values outside this interval is 83.77

hence probability=1/10 =0.1

for sample 2    2 std deviation away values from mean =83.92-/+2*0.0873=83.7454 to 84.0946

values outside this interval are 0

hence probability=0/10 =0.0